# CHESS: Contextual Harnessing for Efficient SQL Synthesis

### v1

@@ 3.4 Preprocessing
To facilitate the information retrieval process outlined in 3.1 efficiently, we preprocess database values and catalogs before executing the pipeline. For database values, we conduct a syntactic search by creating a Locality Sensitive Hashing (LSH) index, as described in entity retrieval. For database catalogs, which contain longer texts requiring semantic understanding, we use a vector database retrieval method to measure semantic similarity.
Locality Sensitive Hashing Indexing of Values.
To optimize the entity retrieval step, we employ a method capable of efficiently searching through large databases, which may contain millions of rows, to retrieve the most similar values. This step doesn’t require perfect accuracy but should retrieve a reasonably small set of similar values, such as a hundred elements. Locality Sensitive Hashing (LSH) is an effective technique for approximate nearest-neighbor searches. It allows us to retrieve database values that are most similar to a given keyword. During preprocessing, we index unique database values using LSH. Then, during the entity retrieval step, we query this index to quickly find the top similar values for a keyword.
@@
As explained in context retrieval, extracting the most semantically relevant pieces of information from database catalogs is crucial for writing a SQL query. These documents can be extensive, with hundreds of pages explaining the entities and their relationships within the database, necessitating an efficient retrieval method. To perform a high-efficiency semantic similarity search, we preprocess the database catalogs into a vector database. During the context retrieval step, we query this vector database to find the most relevant pieces of information for the question at hand. For a more detailed description of our pipeline and the preprocessing phase, we encourage the reader to read Appendix C.
4 Experiments
@@ 4.1 Datasets and Metrics
The Spider dataset (Yu et al., 2018) includes 200 database schemas, with 160 schemas available for training and development, and 40 schemas reserved for testing. Notably, the databases used in the training, development, and test sets are distinct and do not overlap.
@@
The recently introduced BIRD dataset (Li et al., 2024b) features 12,751 unique question-SQL pairs, covering 95 large databases with a combined size of 33.4 GB. This dataset spans 37 professional fields, including sectors such as blockchain, hockey, healthcare, and education. BIRD enhances SQL query generation by incorporating external knowledge and providing a detailed database catalog that includes column and database descriptions, thereby clarifying potential ambiguities. The SQL queries in BIRD are generally more complex than those found in the Spider dataset.
Subsampled Development Set (SDS).
@@
To facilitate ablation studies, reduce costs, and maintain the distribution of the BIRD development set, we subsampled 10% of each database in the development set, resulting in the Subsampled Development Set which we call SDS. This SDS consists of 147 samples: 81 simple, 54 moderate, and 12 challenging questions.
Metrics.
The Spider dataset evaluates SQL queries using two metrics: exact-set-match accuracy (EM) and execution accuracy (EX). Exact-set-match accuracy assesses each clause independently, requiring a perfect match with its corresponding clause in the reference SQL query. A SQL query is deemed correct only if it aligns completely with the reference across all components. However, this metric does not account for data values and has a high false negative rate due to multiple valid SQL formulations for a single question.
Execution accuracy (EX) evaluates the accuracy of the SQL output by comparing the results of the predicted query with those of the reference query when executed on specific database instances. This metric provides a more nuanced understanding of performance by accounting for variations in valid SQL queries that may arise from the same question.
Additionally, the BIRD benchmark introduces the Valid Efficiency Score (VES), which evaluates SQL query performance by considering both accuracy and execution speed. We achieve a similar ranking compared to other models on this metric as well but due to its high variance and dependence on the computational environment we exclude it from the current analysis.
4.2 BIRD Results
As reported in Table 1(a), our approach using proprietary models achieved state-of-the-art execution accuracy on both the development and test sets of BIRD. Our method with open-source LLMs attained the highest performance among all open-source methods. At the time of the submission of this paper, the highest-performing method on the BIRD leaderboard is ExSL + granite-20b-code with an accuracy of 67.86% on the test set. Our approach ranks second with an accuracy of 66.69%.
Table 1: Performance of our proposed method on the BIRD development set and Spider test set, comparing to all published methods.
(a) BIRD results
@@ 4.2 BIRD Results
CHESS + proprietary (ours) 66.69 65.00
@@
(Lee et al., 2024)
CHESS + Open LLMs (ours) 61.5
@@
(Wang et al., 2023)
(b) Spider test set
@@
(Li et al., 2023)
4.3 Spider Results
To evaluate the generalizability of our proposed method beyond the BIRD benchmark, we tested it on the Spider test set without specifically fine-tuning a new model for candidate generation or modifying the in-context learning samples. We followed our default engine setup. The only adjustment we made to our pipeline was the removal of the context retrieval node since the Spider test set lacks column or table descriptions, which are integral to our method. As shown in Table 1(b), our approach achieved an execution accuracy of 87.2% on 2,147 samples from the test set, ranking it as the second-highest performing method among those published. This underscores the robustness of our method across different databases without any modifications. Notably, the best propriety (and undisclosed) method on the Spider test set leaderboard is Miniseek with an accuracy of 91.2%.
@@ 4.4 Ablation Studies
Models Ablation.
Thanks to our efficient retrieval process, which carefully controls the number of tokens passed to LLMs, we can utilize an open-source LLM with a small context window size, specifically Llama-3 with only 8K tokens AI . This contrasts with previous works that predominantly use GPT-4 as their base model (Pourreza and Rafiei, 2024a; Lee et al., 2024; Wang et al., 2023). In Table 2, we present the results of our proposed pipeline using various LLMs from different families on a subsampled development set dataset. The results indicate that our fine-tuned model for candidate generation significantly enhances performance. Notably, Llama-3’s performance surpasses that of GPT-3.5-turbo but does not yet reach the performance levels of GPT-4 in our analysis.
Engine setups EX
(GPT-3.5-turbo, Fine-tuned DeepSeek, GPT-4-turbo) 64.62
(GPT-3.5-turbo, GPT-4-turbo, GPT-4-turbo) 55.78
(GPT-3.5-turbo, GPT-3.5-turbo, GPT-3.5-turbo) 49.65
(Llama-3-70B, Llama-3-70B, Llama-3-70B) 54.42
(Llama-3-70B, Fine-tuned DeepSeek, Llama-3-70B) 59.86
Table 3: The execution accuracy (EX) of the pipeline by removing each component on the (subsampled) dev set.
[…]
Full pipeline 64.62
w/o entity & context retrieval 59.86 -4.76
w/o individual column filtering 61.90 -2.72
w/o table selection 58.50 -6.12
w/o final column selection 59.18 -5.44
w/o revision 57.82 -6.80
with 1-time revision 61.22 -3.40
Modules Ablation.
@@
Table 3 presents the execution accuracy (EX), where different modules or components are omitted. In the configuration without entity and context retrieval, we retrieved a random example and included column descriptions for all columns. This approach highlights the significant impact of our selective retrieval, which outperforms naive context augmentation by 4.76% in execution accuracy. Additionally, we evaluated the effect of removing each submodule within the schema selection module, revealing that table selection is the most critical, contributing a 6.12% increase in performance. The table also illustrates the significant influence of the revision node, with a 6.80% improvement. Increasing the number of revision samples for self-consistency led to higher performance gains, aligning with findings from (Lee et al., 2024).
4.5 Performance Evaluation Across Queries with Varying Complexity
The BIRD benchmark categorizes questions and SQL pairs based on the number and type of SQL keywords used into three classes: easy, moderate, and challenging. In this section, we evaluate the performance of our method 1) using our fine-tuned model for candidate generation, and 2) using GPT-4 without our fine-tuned model. We compare the results to the original GPT-4 baseline as in the BIRD paper, where the question, evidence, and the complete schema with all tables and columns are presented to GPT-4 along with chain-of-thought reasoning prompts.
The analysis is conducted on the SDS dataset, and the results are detailed in Table 4. Our proposed method in both settings significantly improved performance across all classes. This analysis further demonstrates that providing an LLM with all available information can confuse the model and selective retrieval is crucial for achieving higher performance.
Table 4: Comparing the performance of our proposed method in two different settings with naively passing all information to GPT-4 across different difficulty levels on the subsampled development set.
Easy Moderate Challenging Overall
CHESS (with fine-tuning) 65.43 64.81 58.33 64.62
CHESS (w/o fine-tuning) 60.49 50.00 50.00 55.78
GPT-4-turbo (baseline) 54.32 35.18 41.66 46.25
4.6 Evaluation of the Schema Selection
@@
Aside from the ablation studies to measure the effects of context retrieval, schema selection, and revision methods, we also measure their effect on the precision and recall of the tables and columns provided to the model for generating the final SQL. Precision and recall are calculated using the columns and tables that are used in the correct SQL queries as ground truth. As shown in Table 5, each step of the pipeline increases the precision of the selected tables and columns while only slightly decreasing the recall.
By achieving high recall and precision, the model has the most relevant information to generate the correct SQL in a small context window. An example of the reduction in the number of tables and columns of our pipeline is shown in Figure 12 where the final two tables are selected correctly and the final selected five columns include the two correct columns used by the correct SQL.
Table 5: Recall and Precision for Individual Column Filtering, Table Selection, and Column Selection compared to the tables and columns used in the correct SQL.
@@
No Filtering and Selection 1.0 0.33 1.0 0.11
Individual Column Filtering 1.0 0.33 0.98 0.21
Table Selection 0.97 0.89 0.96 0.45
Final Column Selection 0.96 0.90 0.94 0.71
5 Discussion and Limitations
In this paper, we propose a new LLM-powered pipeline, called CHESS, for effective text-to-SQL synthesis. CHESS consists of novel and efficient retrieval and schema pruning methodologies as well as scalable finetuning techniques. Our pipeline achieved state-of-the-art performance among all known methodologies on the challenging BIRD benchmark. Furthermore, we also developed an entirely open-source version of CHESS, that for the first time, surpassed 60% execution accuracy on BIRD, narrowing the performance gap between closed-source and open-source LLMs, while ensuring the privacy of the data.
While our approach increases the capability of text-to-SQL methodologies, the ultimate goal is to bring full automation to the database querying process. For the challenging BIRD dataset, humans still achieve a larger performance on query synthesis and future work should aim to further close this gap. We have quantified the efficacy of each of the components of our pipeline, which presents a guideline for strategies for further enhancements. Each of the information retrieval, schema pruning, and synthesis could benefit from further improvements, although as described in the paper, devising higher precision schema selection methodologies would be a high-impact area for future research that as it stands, might have an outsize impact on end-to-end accuracy.
Appendix A Query Generation with the Correct Context
We also fine-tuned the query generator module with the correct contextual information including only the relevant columns and tables and their descriptions. As shown in 6, the performance with the correct contextual information based on the gold SQL reached 72.4%, underscoring the critical role of retrieving the efficient schema information.
Engine EX
CHESS 64.62
CHESS + correct context 72.4
Appendix B Finetuning the SQL Generator

### v3

@@ 3.2 Preprocessing
To accelerate the IR agent’s tools, retrieve_entity and retrieve_context, and improve their efficiency, we preprocess database values and catalogs before running the system. For database values, we perform a syntactic search by creating a Locality-Sensitive Hashing (LSH) index, as explained in retrieve_entity. For database catalogs, which contain longer texts that require semantic understanding, we use a vector database retrieval method to measure semantic similarity.
Locality Sensitive Hashing Indexing of Values. To optimize the entity retrieval process, we employ a method capable of efficiently searching through large databases, which may contain millions of rows, to retrieve the most similar values. This process doesn’t require perfect accuracy but should return a reasonably small set of similar values, such as around a hundred elements. Locality Sensitive Hashing (LSH) is an effective technique for approximate nearest-neighbor searches, allowing us to retrieve database values most similar to a given keyword. During preprocessing, we index unique database values using LSH. Then, in the retrieve_entity tool, we query this index to quickly find the top similar values for a keyword.
@@
Extracting the most semantically relevant pieces of information from database catalogs is crucial for generating accurate SQL queries. These catalogs can be extensive, with hundreds of pages explaining the entities and their relationships within the database, necessitating an efficient retrieval method. To enable a high-efficiency semantic similarity searches, we preprocess the database catalogs into a vector database. During the context retrieval step, we query this vector database to find the most relevant information for the question at hand.
For a more detailed description of our framework and the preprocessing phase, please refer to Appendix A.
3.3 Adapting Text-to-SQL Systems for Real-World Deployment
We designed our multi-agent framework to allow for flexible combinations of agents, enabling it to adapt to various deployments, database settings, and constraints. In setups where higher computational budgets and more powerful LLMs are available, CHESS can be configured to maximize accuracy by leveraging these resources. Specifically, the Candidate Generator (CG) and Unit Tester (UT) agents can be programmed to generate multiple candidate queries and unit tests, leading to more accurate final SQL queries Brown et al. (2024). This increased sampling allows the system to refine its query generation process, resulting in better performance, as demonstrated in Section 4.2.1.
Interestingly, similar to the findings in Maamari et al. (2024), we observed that when schema sizes are relatively small and the available LLMs are highly capable, the inclusion of the Schema Selector (SS) agent may introduce a precision-recall trade-off that can negatively impact accuracy. In these cases, omitting the SS agent can lead to better results.
However, as we will demonstrate in Section 4.2.2, for larger schemas, even powerful LLMs struggle to maintain performance. In such cases, the collaboration of all four agents is essential for achieving higher accuracy, as the SS agent becomes crucial for managing the complexity of the schema.
Finally, when computational resources are limited or when only weaker LLMs, such as small open-source models, are available to ensure data privacy, CHESS can be reconfigured for efficiency. In these scenarios, we streamline the system by excluding the UT agent and limiting the CG agent to generate only one candidate query. This approach minimizes computational overhead while still synthesizing a reliable SQL query, making CHESS suitable for resource-constrained environments.
4 Experiments and Results
@@ 4.1.1 Datasets
Spider. The Spider dataset (Yu et al., 2018) includes 200 database schemas, with 160 schemas available for training and development, and 40 schemas reserved for testing. Notably, the databases used in the training, development, and test sets are distinct and do not overlap.
@@
Bird. The recently introduced BIRD dataset (Li et al., 2024b) features 12,751 unique question-SQL pairs, covering 95 large databases with a combined size of 33.4 GB. This dataset spans 37 professional fields, including sectors such as blockchain, hockey, healthcare, and education. BIRD enhances SQL query generation by incorporating external knowledge and providing a detailed database catalog that includes column and database descriptions, thereby clarifying potential ambiguities. The SQL queries in BIRD are generally more complex than those found in the Spider dataset.
@@
Subsampled Development Set (SDS). To facilitate ablation studies, reduce costs, and maintain the distribution of the BIRD development set, we subsampled 10% of each database in the development set, resulting in the Subsampled Development Set which we call SDS. This SDS consists of 147 samples: 81 simple, 54 moderate, and 12 challenging questions. For reproducibility, we included the SDS in our GitHub repository.
Synthetic Industrial-scale Database Schema. Production-level text-to-SQL systems often operate on real-world databases with extremely large schemas containing thousands of columns, far exceeding the scope of academic benchmarks like BIRD (Li et al., 2024b). As highlighted in the Lost-in-the-Middle paper (Liu et al., 2024), large language models (LLMs) can struggle with overly long prompts, leading to confusion. However, recent advancements, such as Gemini (Team et al., 2023) and GPT-4o (OpenAI, 2024b), have demonstrated the capability to generate accurate SQL queries without the need to filter prompts or exclude any part of the database schema (Maamari et al., 2024). To assess the scalability of these models on massive databases, we synthesized extremely large datasets to mimic real-world conditions. This involved merging multiple databases from the BIRD training and development sets, resulting in schemas with up to 4,337 columns, the largest in our experiments. This method enabled us to replicate the complexity and challenges of large, production-level databases.
Metrics. In this paper, we primarily used Execution accuracy as the metric for comparison following the previous works (Pourreza et al., 2024; Maamari et al., 2024; Pourreza and Rafiei, 2024a). Execution Accuracy (EX) measures the correctness of the SQL output by comparing the results of the predicted query with those of the reference query when executed on specific database instances. This metric offers a nuanced evaluation by accounting for variations in valid SQL queries that can yield the same results for a given question. Additionally, given that the number of LLM calls and the tokens used in prompts can be substantial in some scenarios, we also report both the total number of tokens and the number of LLM calls required by our approach.
4.2 Different Configurations of CHESS per Deployment Constraints
In this section, we discuss various deployment constraints, detailing how CHESS is configured for each scenario, along with the corresponding experiments and results. First, in section 1, we consider a scenario with high computational budget that allows multiple LLM calls during inference using strong LLMs. For this case, we evaluate the performance of CHESS using the IR, CG, and UT agents, referred to as $\text{CHESS}_{(IR,CG,UT)}$, on the BIRD dataset. Next, in section 4.2.2, we address the challenge of industrial-scale database schemas beyond those in the BIRD dataset, where the reasoning capabilities of powerful proprietary models, such as Gemini-1.5-pro, may be insufficient. To tackle this issue, we introduce the SS agent to the team, demonstrating its effectiveness in handling very large schemas. Then, to accommodate scenarios with limited computational budgets, we exclude the UT agent from the team and restrict the number of candidates generated by CG to one. This configuration, labeled as $\text{CHESS}_{(IR,SS,CG)}$, is evaluated using both less powerful proprietary models and open-source models. Finally, in section 4.2.3, we perform ablation studies to assess the contribution of each component to the end-to-end performance of our system.
4.2.1 $\text{CHESS}_{(IR,CG,UT)}$ for High Computational Budget
Recent studies, such as MCS-SQL (Lee et al., 2024) and CHASE-SQL (Pourreza et al., 2024), have achieved state-of-the-art performance on text-to-SQL benchmarks by scaling LLM calls, generating a large set of candidate SQL queries, and selecting the best candidate among them. Following a similar strategy, we evaluate the performance of our text-to-SQL multi-agent framework by generating 20 candidate SQL queries using the Gemini-1.5-pro model for each sample question. We then create ten natural language unit tests to effectively differentiate correct candidates from incorrect ones. Given that schema linking is unnecessary for small databases within the BIRD benchmark (Maamari et al., 2024), we excluded the SS agent and retained only the IR, CG, and UT agents for this experiment. The performance comparison between our framework with Gemini-1.5-pro model and prior methods is presented in Table 1. As shown in the table, our approach outperforms all previous methods except CHASE-SQL, which utilizes a fine-tuned model specifically designed to select the best candidate SQL query for the BIRD benchmark. In contrast, our framework, without any fine-tuning, surpasses all other methods, demonstrating its robustness across benchmarks without the need for task-specific optimization. Furthermore, we evaluated the performance of the UT agent in two distinct settings. In the first setting, a single LLM call receives one unit test and assigns scores to all candidate queries. In the second setting, a single LLM call receives all unit tests and assigns a score to one candidate query. The results for the first scenario are presented in Table 1, where we achieved a score of 68.31. In contrast, the second scenario yielded a lower score of 66.78, highlighting the significance of allowing the LLM to compare candidate queries when evaluating unit tests.
Table 1: Performance of $\text{CHESS}_{(IR,CG,UT)}$ on BIRD dataset, comparing to the methods with high computation budget.
Method dev EX test EX
CHASE-SQL 73.01 73.00
$\textbf{CHESS}_{(IR,CG,UT)}$ 68.31 71.10
Distillery + GPT-4o 67.21 71.83
MCS-SQL + GPT-4 63.36 65.45
SFT CodeS-15B 58.47 60.37
DTS-SQL + DeepSeek 7B 55.8 60.31
MAC-SQL + GPT-4 59.59 59.59
4.2.2 Introducing the Schema Selection Agent for Large Database Schema
In this section, we use our synthetically generated dataset, the Synthetic Industrial-scale Database Schema, to evaluate the performance of our multi-agent framework on extremely large schemas and analyze the impact of large prompts on the performance of recent LLMs. We assessed the performance of Gemini model using Pass@1 and Pass@5 metrics on BIRD development set using the synthetic dataset as target database. For database sizes smaller than the maximum size (4337), we randomly select columns while ensuring that the ground truth schema is retained. Pass@1 measures the success rate when the model is allowed to generate a single SQL query with execution accuracy of one, while Pass@5 measures the rate of generating at least one correct query from five attempts. We tested the Gemini-1.5-pro model and presented the results in Figure 3, which compares the Pass@1 and Pass@5 performances across varying schema sizes. The results reveal an 11% performance gap between scenarios where the exact ground truth schema (i.e., the correct columns and tables) is provided and those where a large number of columns are included in the prompt. As the number of columns decreases, the model’s performance improves, underscoring the importance of schema linking to filter out irrelevant columns and tables.
Effectiveness of the Schema Selection Agent. To measure the effectiveness of our schema selection agent (SS agent) in dealing with very large schema, we repeated the previous experiment this time including the SS agent. After applying the SS agent to the largest schema with 4,337 columns, the Pass@1 and Pass@5 scores improved to 61% and 63%, respectively, representing a 2% increase compared to the performance without schema linking. This demonstrates the critical role of schema linking in enhancing query generation for very large databases.
Note that the Pass@1 of the system including SS agent has a slightly better performance of the Pass@5 of the system without any schema linking, which shows a higher performance while being more than ×$\times 5$ efficient in terms of used tokens. This suggests that for very large schema the schema selection agent can increase the accuracy as well as the efficiency of the system.
Figure 3: Text-to-SQL Pass@1 and Pass@5 performance of Gemini-1.5-pro model vs. the number of columns in the database.
4.2.3 $\text{CHESS}_{(IR,SS,CG)}$ for Limited Computational Budget
In real-world deployment scenarios, various constraints such as low computational power, limited access to powerful models due to budget restrictions, or data privacy concerns can make the use of high-compute methods with powerful proprietary models (e.g., Lee et al. (2024); Pourreza et al. (2024)) infeasible. To accommodate such cases, we propose a configuration of our multi-agent system, $\text{CHESS}_{(IR,SS,CG)}$, that achieves high performance while adhering to these constraints.
To limit the number of LLM calls, we exclude the UT agent from the framework and restrict the number of candidate generations and revisions by GC to only one and three times, respectively. Because we cannot rely on multiple LLM calls to generate candidate queries, it becomes crucial to be more efficient with the schema passed to the candidate generator. Therefore, the Schema Selector (SS) plays a vital role in enhancing efficiency by narrowing down the schema to the most relevant subset before candidate generation. Moreover, to demonstrate the performance of the system using less powerful models, we conduct experiments using older proprietary models such as GPT-3.5/4-turbo, as well as open-source models like Llama-3-70B and a fine-tuned DeepSeek model.
BIRD Results. Since the test set of the BIRD benchmark is not available, we conducted our ablations and performance evaluations on the development set. We assessed our proposed method using both 1) proprietary and 2) open-source models. In the first case, we utilized a fine-tuned DeepSeek Coder model for candidate generation, GPT-3.5-turbo for column filtering, and GPT-4-turbo for the remaining LLM calls. In the second case, we used our fine-tuned DeepSeek Coder model for candidate generation, with all other LLM calls handled by Llama-3-70B.
As reported in Table 2, our approach using proprietary models achieved a high execution accuracy on both the development and test sets of BIRD. We want to highlight the efficiency of our method in this configuration where we only call GPT-4-turbo model 6 times where some other methods such as MCS-SQL performs around 100 LLM calls.
CHESS using Open-source Models. In most of the industrial use cases, the privacy of the data is at the highest priority, which limits the use of proprietary models as in Pourreza et al. (2024), Maamari et al. (2024), and Lee et al. (2024). For such cases, companies would prefer to deploy a text-to-SQL system using only open-source models completely on-premis. Our method with open-source LLMs attained the highest performance among all open-source methods, making it very suitable for these scnearios.
Table 2: Performance of $\text{CHESS}_{(IR,SS,CG)}$ on the BIRD dataset, comparing to the methods with low computation budget.
@@ 4.2.3 $\text{CHESS}_{(IR,SS,CG)}$ for Limited Computational Budget
$\textbf{CHESS}_{(IR,SS,CG)}$ + proprietary 66.69 65.00
@@
$\textbf{CHESS}_{(IR,SS,CG)}$ + Open LLMs 61.5
@@
Table 3: Performance of $\text{CHESS}_{(IR,SS,CG)}$ on the Spider test set, comparing to all published methods.
@@
Spider Results. To evaluate the generalizability of our proposed method beyond the BIRD benchmark, we tested it on the Spider test set without specifically fine-tuning a new model for candidate generation or modifying the in-context learning samples. We followed our default engine setup. The only adjustment we made to our setting was the removal of the retrieve_context tool since the Spider test set lacks column or table descriptions, which are integral to our method. As shown in Table 3, our approach achieved an execution accuracy of 87.2% on the test set, ranking it as the second-highest performing method among those published. This underscores the robustness of our method across different databases without any modifications. Notably, the best propriety (and undisclosed) method on the Spider test set leaderboard is Miniseek with an accuracy of 91.2%.
Table 4: The execution accuracy (EX) of the framework by removing each tool on the (subsampled) development set.
Tools EX $\Delta$EX
All tools 64.62
w/o retrieve_entity & _context 59.86 -4.76
w/o filter_column 61.90 -2.72
w/o select_tables 58.50 -6.12
w/o select_columns 59.18 -5.44
w/o revise 57.82 -6.80
with 1-time revise 61.22 -3.40
@@ 4.3 Ablation Studies
Table 4 presents the execution accuracy (EX), where different modules or components are omitted. In the configuration without entity and context retrieval, we retrieved a random example and included column descriptions for all columns. This approach highlights the significant impact of our selective retrieval, which outperforms naive context augmentation by 4.76% in execution accuracy. Additionally, we evaluated the effect of removing each tool of the SS agent, revealing that the select_tables tool is the most critical, contributing a 6.12% increase in performance. The table also illustrates the significant influence of the revision tool, with a 6.80% improvement. Increasing the number of revision samples for self-consistency led to higher performance gains, aligning with findings from (Lee et al., 2024).
Unit Tests Improve Performance. To assess the effectiveness of the UT agent in generating high-quality unit tests for distinguishing between correct and incorrect SQL candidates, we evaluated the performance of our agent-based framework by varying the number of unit tests and select the candidate with the highest score from the 20 candidates. Figure 4 illustrates the performance across different numbers of unit tests. As shown in the figure, performance improves as the number of unit tests increases, peaking at 10 tests. Beyond this point, the performance plateaus, indicating that additional tests provide diminishing returns.
Figure 4: Text-to-SQL performance of CHESS with respect to different number of unit tests on BIRD development set.
@@
Evaluation of the Schema Selection Agent In addition to conducting ablation studies to assess the impacts of context retrieval, schema selection, and revision methods, we evaluate their specific influence on the precision and recall of tables and columns identified by the model for generating final SQL queries. Precision and recall are calculated by comparing the tables and columns used in correct SQL queries as ground truth. As shown in Table 5, the schema selection agent improves the precision of selected tables and columns, with only a slight reduction in recall.
This balance of high precision and recall enables the model to work with the most relevant information in a compact context window, enhancing the accuracy of SQL generation. Figure 16 illustrates this process, where the schema selection (SS) agent successfully narrows down the tables and columns, selecting two tables and five columns, including the two columns used in the correct SQL query. For a comprehensive example of the schema selection process, please refer to Appendix E.
Furthermore, Figure 5 illustrates how the schema selection agent effectively narrows down the number of columns at each step, adapting the remaining columns based on the complexity of the question rather than the database schema itself. This approach consistently reduces the columns to around 10, making our method scalable and applicable across databases of varying complexity by focusing only on the most relevant information for SQL generation.
Table 5: Precision and Recall of schema items for tables and columns used in correct SQL, evaluated after applying each tool by the schema selection agent: filter_columns, select_tables, and select_columns.
@@
No Filtering or Selection 1.0 0.33 1.0 0.11
filter_columns 1.0 0.33 0.98 0.21
select_tables 0.97 0.89 0.96 0.45
select_columns 0.96 0.90 0.94 0.71
Figure 5: The average number of columns remaining after each stage of the schema selection process across 11 BIRD dev databases. The initial count shows the total columns in each database, and the schema selection agent consistently reduces this number to around 10 columns for candidate generation, regardless of database complexity.
Performance Evaluation Across Queries with Varying Complexity The BIRD benchmark categorizes questions and SQL pairs based on the number and type of SQL keywords used into three classes: easy, moderate, and challenging. In this section, we evaluate the performance of our method 1) without computation constraints by using the IR, CG and UT agents, and 2) with limited computational budget by using the IR, SS, and CG agents. We compare the results to the original GPT-4 baseline as in the BIRD paper, where the question, evidence, and the complete schema with all tables and columns are presented to GPT-4 along with chain-of-thought reasoning prompts.
The analysis is conducted on the SDS dataset, and the results are detailed in Table 6. Our proposed method significantly improves performance across all classes in both settings. This demonstrates the effectiveness of our multi-agent framework in enhancing performance across varying difficulty levels and constraints.
Table 6: Comparing the performance of CHESS in two different settings with a naive GPT-4 baseline where we pass all the information in its context, across different question difficulty levels.
Easy Moderate Challenging Overall
$\textbf{CHESS}_{(IR,CG,UT)}$ 73.48 62.99 59.31 68.31
$\textbf{CHESS}_{(IR,SS,CG)}$ 65.43 64.81 58.33 64.62
GPT-4-turbo (baseline) 54.32 35.18 41.66 46.25
5 Conclusion and Limitations
In this paper, we introduced CHESS, a multi-agent framework capable of handling complex, industrial-scale databases. CHESS devises 4 specialized agents, namely, Information Retriever, Schema Selector, Candidate Generator, and Unit Tester, to tackle key challenges such as efficient data retrieval, schema pruning, SQL query generation, and validation. We provide extensive ablation studies to show the effectiveness of each of these agents and their tools. We demonstrate that for large industrial-scale schema sizes, methods solely relying on scaling sampling at inference time become less effective, whereas CHESS’s schema selection agent can deliver a high accuracy while processing less than ×$\times 5$ tokens. We also show that CHESS achieves state-of-the-art performance when controlled for computational budget. CHESS achieves 71.10% accuracy on the challenging BIRD test set, within 2% of the leading method (a proprietary approach), while requiring approximately 83% fewer LLM calls.
As discussed in the experiments section, there is a significant need for an extensive text-to-SQL benchmark that accurately reflects the challenges posed by large-scale databases. In our paper, we attempted to achieve this by combining different databases; however, having access to even larger real-world databases can meaningfully advance research in this domain. Additionally, previous studies (Pourreza et al., 2024) have shown that finetuning can improve LLM performance in query selection. As future work, we plan to finetune models specifically for test-case generation and evaluation to further improve the performance.
Appendix A Implementation Details
A.1 Locality Sensitive Hashing Indexing of Database Values
Our goal in the retrieve_entity tool is to retrieve database values that most closely match a set of keywords derived from the question. It is important to recognize that keywords from the question may not exactly correspond to the database values due to potential typos, variations in expression, or the common scenario where users are unaware of the precise format used to store data in the database. This reality demands a retrieval strategy that is both robust and adaptable, capable of accommodating such discrepancies. Relying solely on exact match retrieval, as suggested in prior studies [Li et al., 2024a], may not be sufficiently effective.
Locality Sensitive Hashing (LSH) is an efficient technique for approximate nearest neighbor searches, which allows us to retrieve the most similar values to a keyword in the database. In the pre-processing stage, we index unique values in the database using LSH. Then, in the retrieve_entity tool of the IR agent, we query this index to rapidly find the top similar values to a keyword. Our approach involves using LSH queries to retrieve the top 10 similar values, after which we compute the edit distance and semantic similarity between the keyword and these values to further refine the results.
To simultaneously utilize edit distance and embedding similarity, we first identify the top 10 values closest to each keyword based on cosine similarity between their embedding vectors (obtained using OpenAI text-embedding-3-small OpenAI [2024a]) and the keyword’s embedding vector. We then filter out values that fall below a specific threshold. Finally, for each keyword and column, we retain only the value that has the smallest edit distance.
We observed a significant reduction in time complexity, from 5 minutes to 5 seconds, using this method compared to a naive approach of computing the edit distance for all unique values in the database on the fly. While computing edit distance is proportional to the size of the database, significantly increasing the time complexity for processing a single question, using LSH allows us to index values in the pre-processing step and, during entity retrieval, rapidly query the index to find the most similar values to a keyword in a much more time-efficient manner.
A.2 Vector database
Each database schema in the BIRD benchmark Li et al. [2024b] includes detailed descriptions for columns, specifying the contents of each column and the values for categorical columns. Providing these descriptions to the model is essential for guiding the SQL query generation process. However, incorporating all descriptions in the prompt for the weaker open-source models can overwhelm the model, potentially leading to the generation of incorrect SQL queries, as observed in section 5. It is important to note that the database catalog in the BIRD benchmark provides a relatively limited view of database metadata. In contrast, real-world production-level databases often contain more diverse information, including value ranges, constraints, and usage instructions for each table. Our proposed method can effectively utilize this extensive metadata to enhance performance.
To evaluate the relevance of descriptions to a given question, we employ embedding similarity [OpenAI, 2024a], which quantifies the semantic similarity between the question and each description. To enhance the efficiency of the retrieval process, we pre-process the descriptions and created the embedding vectors for each of them and stored in a vector database, utilizing ChromaDB in our implementation. For the retrieve_context tool of the IR agent, we query this vector database to identify descriptions that are most semantically aligned with the question. This targeted approach ensures that only the most pertinent information is provided to the model, thereby improving the accuracy of the generated SQL queries.
A.3 Local Column filtering
In the filter_column tool, decisions to retain a column for subsequent steps are made independently, without considering relative information. The data provided to the LLM for column filtering includes: 1) the table name, 2) the column name, 3) the data type, 4) descriptions, if retrieved by the IR agent, and 5) database values, if retrieved by retrieve_entity tool. To enhance the model performance and make the task definition clear to the model, we used few-shot samples for this tool.
Some key columns for SQL generation, which we call linking columns, such as those with foreign and primary key constraints, are crucial for writing SQL queries. For instance, questions about counting entities often requires primary keys, and joining tables necessitates foreign key columns. However, in the filter_column, select_table select_column tools, some of these essential columns may be initially rejected because they do not semantically relate to the given question. Despite this, these columns are indispensable for SQL generation. Therefore, in all of our sub-modules, we consistently retain foreign key and primary key columns, irrespective of the outputs from column selection and filtering processes.
A.4 Revise Tool
Revising generated candidate SQL queries is a critical aspect for the Candidate Generator Agent. In addition to the database schema, the question, and the candidate SQL query, we also provide the execution result of the SQL query. This gives the LLM an opportunity to view the retrieved data and revise the SQL query accordingly. This process mirrors human behavior when writing complex SQL queries; typically, we start with a draft query and refine it based on the results of its execution. Furthermore, this method allows the LLM to make necessary adjustments to the SQL query in instances of execution syntax errors.
In this step, we also incorporate instructions derived from our error analysis F to guide the model towards generating correct SQL queries. For instance, as shown in 6, we guided the model to ensure that all requested columns are included in the SQL query. In this specific example, the revise tool identified a missing column and successfully added it to the query.
Figure 6: An example of the revise tool to fix missing columns in a candidate query.
Appendix B Finetuning the Model for Generate_Candidate_Query